Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
bersanetti, pinto, ruggi - 0:39 Thursday 08 April 2021 (51349) Print this report
ISC shift: DRMI and Green lock recovery

Today's shift was dedicated to continue the recovery of the green beam alignment in the arms, started in the morning shift. We started the alignment of the West arm, working on the setpoints of the Galvos. The procedure appeared to be quite difficult, since it was possible to obtain enough power on the ALS_WEB/NEB_PD_REFL signals, while the power on the transmission signals was poor. We eventually found a good enough configuration (good setpoints for the Galvos); around 18:50:00 UTC we were able to obtain a stable lock on the west arm, and allow the Drift Control to compensate the remaining part of misalignment. At this point, with the same procedure, we started to align the North arm, which at first showed the same issues of the West regarding the Transmission power, but at 20.20 UTC we found a good alignment allowing us to lock and to engage the Drift Control to compensate just as it was done for the West arm. For reference, in Fig.1 is reported the trend of the ALS_WEB/NEB_PD_REFL_LP and ALS_CEB_WARM/NARM_BEAT_DC signals for the duration of the entire shift.

A more fine alignment needs to be continued in the following shifts, in order to further increase the Transmission power, together with the improvement of the new alignment strategy started in the morning shift, using the new lines BPC_SCAN_FAST and BPC_SCAN_SLOW.

At 21.00.00 UTC we started to work on the CITF in standalone, with the end mirrors misaligned. We relocked it quite easily on the 1f signals, but we immediately discovered that the power of the B4_112MHz_mag signal was much lower than usual (about a factor 10); at the same time we found that with the usual gain the MICH loop UGF was very high (around 75 Hz), a reduction of the MICH loop gain cured the latter, although the cause must be investigated, while the sidebands issue remains unsolved. We misaligned the SR mirror in-lock, and the sideband power increased a bit, but way less than needed, however, this was clearly a misaligned position as unlocking made well visible the secondary reflections on the B1p camera. All of this will be investigated.

About the PR alignment, we could close it in full bandwidth both on TX and TY, without exciting the 11.5 Hz noise, so the vertical correction fix is still to be tested.

In the end, we realigned and relocked the cavities on the IR beam.

Images attached to this report
Comments to this report:
bersanetti - 2:26 Thursday 08 April 2021 (51350) Print this report

About the DRMI, what really happened is explained in the attached Figure:

  • during the lock acquisition, done with the DRMI_LOCK Automation node, we use, as we always do since we lock the DRMI successfully, the DRMI_TRIGGER_IN2 combination of triggers (B4_12MHz_mag high, B4_112MHz_mag high, B1p_DC low);
  • however, inside the DRMI_TRIGGER_IN main trigger it was still loaded the _IN3 combination, which does not use a high threshold on B4_112MHz_mag; therefore, the lock engaged in a bad place for SRCL which, given the loose nature of the loop, somehow locked in some way, but with very small values on B4_112MHz_mag; once the SR mirror was misaligned, this caused an increase of the B4_112MHz_mag signal, making the _IN2 to finally trigger and the main trigger went to 2;
  • this was a leftover from manual actions done at 07h51m32 UTC, and never reverted or communicated;
  • once the DRMI_LOCK node was used to unlock and go to DOWN, everything was reset correctly and the following lock was acquired with the correct trigger, although the mix of a maybe too low threshold and the still misaligned position of SR (for testing) somehow hid this effect and we could not see this on the spot;
  • the lock of the DRMI will be re-checked tomorrow.

Reminders:

  • manual actions must be reverted or communicated;
  • we have an automation for the DRMI.
Images attached to this comment
Search Help
×

Warning

×